OpenStudy (anonymous):

how many solutions does −4+5(x+1)=2x+3(x+1) have

OpenStudy (anonymous):

lhd id -4+5x+5=5x+1 and rhs is 2x+3(x+1)=5x+3 from here we have 5x+1=5x+3 that is 1=3, which is impossible hence no solution

OpenStudy (anonymous):

@oqijh3rli